OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) The product of two positive integers is 98. if they differ by 7, what are the integers?

OpenStudy (anonymous):

7 and 14

OpenStudy (saifoo.khan):

xy = 98 x - y = 7

OpenStudy (saifoo.khan):

two equations!

OpenStudy (anonymous):

can you show the solutions?

OpenStudy (anonymous):

xy = 98 x - y = 7 => (x -y)^2 = (x+y)^2 - 4xy =49 => (x+y) = 21 => x =14 , y =7